# Olaph: Improving Factuality in Biomedical Long-form Question Answering

### v1

@@ Abstract
In the medical domain, numerous scenarios necessitate the long-form generation ability of large language models (LLMs). Specifically, when addressing patients’ questions, it is essential that the model’s response conveys factual claims, highlighting the need for an automated method to evaluate those claims. Thus, we introduce MedLFQA, a benchmark dataset reconstructed using long-form question-answering datasets related to the biomedical domain. We use MedLFQA to facilitate the automatic evaluations of factuality. We also propose Olaph, a simple and novel framework that enables the improvement of factuality through automatic evaluations. The Olaph framework iteratively trains LLMs to mitigate hallucinations using sampling predictions and preference optimization. In other words, we iteratively set the highest-scoring response as a preferred response derived from sampling predictions and train LLMs to align with the preferred response that improves factuality. We highlight that, even on evaluation metrics not used during training, LLMs trained with our Olaph framework demonstrate significant performance improvement in factuality. Our findings reveal that a 7B LLM trained with our Olaph framework can provide long answers comparable to the medical experts’ answers in terms of factuality. We believe that our work could shed light on gauging the long-text generation ability of LLMs in the medical domain. Our code and datasets are available at https://github.com/dmis-lab/OLAPH.
@@ 1 Introduction
With the increasing versatility and exceptional performance of large language models (LLMs), their utilization in the medical or clinical domain is expanding rapidly [1, 2, 3, 4, 5, 6, 7]. One of the greatest advantages of LLMs in these domains is their capability to assist or even replace physicians’ tasks [8, 9]. This includes scenarios such as question answering (multi-choice [10, 11, 12, 13, 14] or span-based [15]), reporting a patient’s Electronic Health Record [3, 16], and conversations based on patient inquiries [17]. In the medical domain, numerous situations necessitate the long-form text-generation ability of LLMs. Among these, answering questions posed by patients demands conveying factual and crucial claims, highlighting the necessity for an automated method to evaluate these responses.
@@
To address this challenge, it is important to measure the ability of open-foundation LLMs to answer in a long-form text. Thus, we aim to verify it through long-form question-answering (LFQA) tasks. LFQA is a task that requires elaborate and in-depth answers to open-ended questions [18, 19]. Here, two main challenging points arise: One is that models should not hallucinate or generate false information [20, 21, 22]. For example, when a patient asks, what could be causing the white tongue? the response should convey crucial information about why white tongue occurs and its causes (e.g., white tongue is usually caused by a buildup of bacteria and dead cells on the surface of the tongue) while ensuring that incorrect information (e.g., it is usually harmful and permanent) is not provided.
@@
Another challenge lies in the difficulty of automatically evaluating long-text responses. Existing tasks such as summarization or LFQA assess whether appropriate words are used and the semantic meaning is well encapsulated [20, 23, 24, 25, 26]. Furthermore, other methods consist of manually verifying the responses generated by LLMs using human annotators to ensure high factuality and absence of hallucination which are cost-ineffective and labor-intensive [27, 28, 29]. In particular, in the medical field, it’s also important to ensure that the information provided is accurate, up-to-date, and comprehensible to practitioners and patients alike. Developing reliable automatic evaluation methods would greatly enhance the efficiency and scalability of these assessments, leading to rapid and extensive advancements in the research field by reducing reliance on human evaluators.
@@
To this end, we aim to gather existing LFQA datasets and reconstruct them as a benchmark for the automatic evaluation of medical responses. MedLFQA allows evaluating an LLM’s response in detail: whether they effectively include the keywords necessary to answer the question, whether they are semantically similar to the answer, and whether they accurately include crucial claims without delivering hallucinated information. Furthermore, we employ GPT-4 [30] to generate long-form answers and statements if needed (Section 3.1). For validation, we assess the answers and statements through three medical experts for pairwise evaluation. Thus, we identify that GPT-4 generated responses are reliable to use as the LFQA benchmark datasets (Section 3.2).
@@
We then introduce a simple and novel framework Olaph (Optimizing Large language models’ Answers with Preferences of mitigating Hallucination), which leverages automatic evaluation to generate synthetic preference sets that can help align the model with preferred responses. Our Olaph framework is composed of iterative learning through preference optimization on the synthetic preference sets. We first leverage supervised fine-tuning (SFT) to tailor a pre-trained LLM to a question-answering task [31] (Section 4.1). Then, we derive $k$ sampled predictions using temperature scaling [32] (Section 4.2). We iteratively train direct preference optimization [33] by selecting the response with the highest score based on fine-grained evaluation metrics among sampling-based predictions (Section 4.3). Overall, our framework generates synthetic preference sets using automatic evaluation metrics and iteratively trains LLMs with preferred responses the model generates.
@@
Our findings reveal that learning through our Olaph framework step-by-step can enhance long-text generation abilities by prioritizing factuality, semantic similarities, and word composition. We find that making a preference set with self-generated responses based on a wide range of evaluation criteria and iteratively training on the set increases the desired abilities in a long-text generation. Our findings also highlight that, even on evaluation metrics not used during training, LLMs equipped with our Olaph framework demonstrate significant performance improvement in factuality. Surprisingly, 7B models trained with our framework can generate long-form answers comparable to medical experts’ answers which are proven to be high-quality answers.
@@
Overall, our contributions are as follows: (1) We introduce MedLFQA, a benchmark dataset with restructured formats of current biomedical LFQA benchmark datasets that enables automatic evaluation of the long-text generation ability of open foundation LLMs. (2) In this process, we constitute two statements that can evaluate factuality through medical claims originated by long answers, aiding in a comprehensive understanding of long-text generation abilities. (3) We introduce the simple and novel Olaph framework, which leverages automatic evaluation to generate synthetic preference sets and employs iterative learning through preference optimization. (4) In our findings, we demonstrate that 7B models can generate long answers comparable to the medical experts’ answers in terms of factuality.
@@
LiveQA [34] (Q, A) → (Q, A, MH, NH) 100 4 82.8 2.9 3.0
MedicationQA [35] (Q, A) → (Q, A, MH, NH) 666 24 55.5 2.6 2.3
HealthSearchQA [1] (Q) → (Q, A, MH, NH) 3,077 96 118.8 2.6 2.3
K-QA Golden [22] (Q, A, MH, NH) 201 1 88.5 4.4 3.5
K-QA Silver [22] (Q) → (Q, A, MH, NH) 904 106 99.9 2.4 2.0
@@ 2.1 Long-Form Question Answering
Long-form question answering (LFQA) is a task requiring elaborate and in-depth answers to open-ended questions [18, 19, 36]. In the biomedical and clinical domains, numerous scenarios necessitate the long-text generation ability of LLMs. Among these, we focus on the situation in answering questions posed by patients which requires to evaluate the answer comprehensively. We thus adopt biomedical LFQA benchmark datasets to explore how well open foundation LLMs respond to the content that consumers or patients typically inquire about, utilizing benchmarks that gather such inquiries [1, 22, 34, 35].
@@ 2.2 Evaluating Long-Text Generation
The main challenge in conducting comprehensive research on the LFQA benchmark is the difficulty in automatic evaluation [37]. Prior works provide various metrics for evaluating language models’ responses such as comparing the quality of machine-generated text to reference text [38, 39] and capturing non-trivial semantic similarities [40, 41, 42]. With the increasing demand for using responses generated by LLMs, concurrent research also focuses on whether these responses accurately contain factual content and avoid generating false knowledge (i.e., hallucination) [21, 43, 44, 45, 46, 47, 48, 49, 50, 51].
A widely known metric that can be used to measure factuality is FActScore [20], which decomposes LLM responses into atomic facts and checks if they are supported by the source text. Additionally, there are metrics like Hallucination and Comprehensiveness [22] that measure the inclusion of crucial claims in the clinical domain. In detail, Hallucination [22] is a metric used to measure how many clinical claims are contradicted by the response of language models ($\hat{P}$). We compute the score as below,
@@
where $S$ refers to all statements containing Must Have (MH) and Nice to Have (NH) statements (i.e., $|S|=|MH|+|NH|$). Also, Comprehensiveness [22] is a metric used to measure how many clinically crucial claims are included in the response of language models. We compute the score as follows:
@@
To predict the entailment of the response, we use a classification model BioBERT [52] trained on NLI datasets [53, 54] on behalf of GPT-3.5-turbo due to the costs of API Calls. We will describe the usage of these statements in the following section (Section 3.1). Our work is based on using these fine-grained and diverse evaluation metrics to understand how LLMs generate long-form text prioritizing factuality, semantic similarities, and word composition.
Figure 1: Current LFQA benchmark datasets lack comprehensive evaluation criteria, featuring just a pair of questions and answers (or not even an answer). In MedLFQA, we provide GPT-4 generated answers as well as two crucial statements to address this limitation. For instance, a well-generated GPT-4 response provides information on the definition, advantages, disadvantages, and side effects of Lexapro in response to a patient’s inquiry about it. Additionally, the answers and statements are structured to enable assessment of how closely the LLM response aligns with the correct answer in terms of factuality, semantic similarity, and word composition.
@@ 3.1 Reconstruction of Biomedical Long-Form Question-Answering Datasets
The essential part of answering the patient’s question is conveying factual claims without false knowledge. To this end, the authors from [22] provide 1,212 patient questions originating from real-world conversations held on AI-driven clinical platform (i.e., K Health) containing long-form answers and two optional statements: Must Have Statements indicating that a model must include this statement to be medically accurate (e.g., providing all contraindications for a drug) and Nice to Have Statements indicating that the statements are supplemental (e.g., providing additional conditions where this drug may be helpful). These two statements provide an effective way to conduct an automatic evaluation of identifying factuality. Although the pairs of questions and answers are curated by medical experts, the dataset containing long-form answers is only limited to 202 pairs.
In this work, we introduce MedLFQA, which is constructed by expanding and reformulating current LFQA benchmark datasets to evaluate models’ responses automatically. To this end, we gather four biomedical LFQA datasets: LiveQA [34], MedicationQA [35], HealthSearchQA [1], and K-QA [22]. We describe the statistics of the benchmark datasets in Table 1. Each MedLFQA instance is comprised of four components: question (Q), long-form answer (A), Must Have statements (MH), and Nice to Have statements (NH). Specifically, LiveQA and MedicationQA datasets contain patients’ questions and their medical experts’ answers. HealthSearchQA only includes patients’ questions without their answers and crucial claims. In the K-QA dataset, the remaining examples (83%) that only consist of consumer questions are referred to as the K-QA Silver dataset.
In detail, if a medical expert’s answer exists, we create the two statements by decomposing the answer. For datasets containing only patients’ questions, we generate answers and statements using proprietary large language models such as GPT-4.111We provide detailed prompt in Appendix Table 7 For example, Figure 1 shows that the long-form answer generated by GPT-4 contains essential information, such as the pros and cons effects of Lexapro, compared to the golden answer that is curated with medical experts. We qualify the generated answers and statements through medical experts and provide the details in further section 3.2.
@@ 3.2 Qualification of Generated Answers and Statements
Here, we qualify the validity of predictions generated by GPT-4 using the K-QA golden dataset, whose answers are curated by medical experts. In order to assess a better response, we employ nine evaluation criteria from MedPALM: alignment with medical consensus (MC), reading comprehension (RC), knowledge recall (KC), reasoning (R), inclusion of irrelevant content (IRC), omission of important information (OII), potential for demographic bias (PDB), possible harm extent (PHE), possible harm likelihood (PHL). [^2] We describe the details of these evaluation criteria in Appendix A.1 Using the criteria, we conduct a pairwise evaluation between GPT-4 predictions and K-QA golden answers through medical experts. [^3] We employ three residents of Seoul National University Hospital (https://www.snuh.org/intro.do)
@@
We note that GPT-4 generated answers could be erroneous and hallucinated with false knowledge. However, we believe that GPT-4 reached the plateau of solving medical problems even better than medical experts in some instances [1, 55, 56, 57], which could be interpreted that our work will have promising directions for future long-text generation in the medical domain. We hope that correcting false knowledge or wrong answers instead of pairwise evaluation could help the advancement of the research field more rapidly and extensively.

### v2

@@ Abstract
[…]
@@ 1 Introduction
With the increasing versatility and exceptional performance of large language models (LLMs), their utilization in the medical or clinical domain is expanding rapidly (Singhal et al., 2023; Chen et al., 2023; Thirunavukarasu et al., 2023; Sun et al., 2024; Tu et al., 2024; Labrak et al., 2024; Jeong et al., 2024). One of the greatest advantages of LLMs in these domains is their capability to assist or even replace physicians’ tasks (Egli, 2023; Tian et al., 2024). This includes scenarios such as question answering (multi-choice (Jin et al., 2021; Hendrycks et al., 2020; Jin et al., 2019; Pal et al., 2022; Xiong et al., 2024) or span-based (Krithara et al., 2023)), reporting a patient’s Electronic Health Record (Thirunavukarasu et al., 2023; Yang et al., 2022), and conversations based on patient inquiries (Liévin et al., 2024). In the medical domain, numerous situations necessitate the long-form text-generation ability of LLMs. Among these, answering questions posed by patients demands conveying factual and crucial claims, highlighting the necessity for an automated method to evaluate these responses.
@@
To address this challenge, it is important to measure the ability of open-foundation LLMs to answer in a long-form text. Thus, we aim to verify it through long-form question-answering (LFQA) tasks. LFQA is a task that requires elaborate and in-depth answers to open-ended questions (Fan et al., 2019; Stelmakh et al., 2022). Here, two main challenging points arise: One is that models should not hallucinate or generate false information (Min et al., 2023; Wei et al., 2024; Manes et al., 2024). For example, when a patient asks, what could be causing the white tongue? the response should convey crucial information about why white tongue occurs and its causes (e.g., white tongue is usually caused by a buildup of bacteria and dead cells on the surface of the tongue) while ensuring that incorrect information (e.g., it is usually harmful and permanent) is not provided.
@@
Another challenge lies in the difficulty of automatically evaluating long-text responses. Existing tasks such as summarization or LFQA assess whether appropriate words are used and the semantic meaning is well encapsulated (Min et al., 2023; Falke et al., 2019; Laban et al., 2022; Fabbri et al., 2022; Krishna et al., 2023). Furthermore, other methods consist of manually verifying the responses generated by LLMs using human annotators to ensure high factuality and absence of hallucination which are cost-ineffective and labor-intensive (Liu et al., 2023b; Fu et al., 2023; Liu et al., 2023a). In particular, in the medical field, it’s also important to ensure that the information provided is accurate, up-to-date, and comprehensible to practitioners and patients alike. Developing reliable automatic evaluation methods would greatly enhance the efficiency and scalability of these assessments, leading to rapid and extensive advancements in the research field by reducing reliance on human evaluators.
@@
To this end, we aim to gather existing LFQA datasets and reconstruct them as a benchmark for the automatic evaluation of medical responses. MedLFQA allows evaluating an LLM’s response in detail: whether they effectively include the keywords necessary to answer the question, whether they are semantically similar to the answer, and whether they accurately include crucial claims without delivering hallucinated information. Furthermore, we employ GPT-4 (OpenAI, 2023b) to generate long-form answers and statements if needed (Section 3.1). For validation, we assess the answers and statements through three medical experts for pairwise evaluation. Thus, we identify that GPT-4 generated responses are reliable to use as the MedLFQA benchmark (Section 3.2).
@@
We then introduce a simple and novel framework Olaph (Optimizing Large language models’ Answers with Preferences of mitigating Hallucination), which leverages cost-effective and automatic evaluation to generate synthetic preference sets that can help align the model with preferred responses. Our Olaph framework is composed of iterative learning through preference optimization on the synthetic preference sets. We first leverage supervised fine-tuning (SFT) to tailor a pre-trained LLM to a question-answering task (Ouyang et al., 2022) (Section 4.1). Then, we derive $k$ sampled predictions using temperature sampling (Guo et al., 2017) to construct synthetic preference set by adopting cost-effective and multifaceted automatic evaluations (Section 4.2). Then, we construct a preference set in every steps using previous-step models with self-generated responses and iteratively train with alignment tuning until convergence (Section 4.3 and 4.4). Overall, our framework generates synthetic preference sets using automatic evaluation metrics and iteratively trains LLMs with preferred responses the model generates.
@@
Our findings reveal that learning through our Olaph framework step-by-step can enhance long-text generation abilities by prioritizing factuality, semantic similarities, and word composition. We find that making a synthetic preference set with self-generated responses based on a wide range of evaluation criteria and iteratively training on the set increases the desired abilities in a long-text generation. Our findings also highlight that, even on evaluation metrics not used during training, LLMs equipped with our Olaph framework demonstrate significant performance improvement in factuality. Surprisingly, 7B models trained with our framework can generate long-form answers comparable to medical experts’ answers which are proven to be high-quality answers.
@@
Overall, our contributions are as follows: (1) We introduce MedLFQA, a benchmark dataset with restructured formats of current biomedical LFQA benchmark datasets that enables automatic evaluation of the long-text generation ability of open foundation LLMs. (2) In this process, we constitute two statements that can automatically evaluate factuality cost-effectively through medical claims originated by long answers, aiding in a comprehensive understanding of long-text generation abilities. (3) We introduce the simple and novel Olaph framework, which leverages automatic evaluation to generate synthetic preference sets and employs iterative learning through preference optimization. (4) In our findings, we demonstrate that 7B models can generate long answers comparable to the medical experts’ answers in terms of factuality.
@@
LiveQA (Abacha et al., 2017) (Q, A) → (Q, A, MH, NH) 100 4 82.8 2.9 3.0
MedicationQA (Abacha et al., 2019) (Q, A) → (Q, A, MH, NH) 666 24 55.5 2.6 2.3
HealthSearchQA (Singhal et al., 2023) (Q) → (Q, A, MH, NH) 3,077 96 118.8 2.6 2.3
K-QA Golden (Manes et al., 2024) (Q, A, MH, NH) 201 1 88.5 4.4 3.5
K-QA Silver (Manes et al., 2024) (Q) → (Q, A, MH, NH) 904 106 99.9 2.4 2.0
@@ 2.1 Long-Form Question Answering
Long-form question answering (LFQA) is a task requiring elaborate and in-depth answers to open-ended questions (Fan et al., 2019; Stelmakh et al., 2022; Krishna et al., 2021). In the biomedical and clinical domains, LFQA is essential for effectively integrating AI into real-world applications. Despite its importance, there has been relatively little effort to construct patient-centered LFQA datasets due to its domain specificity. In other words, numerous scenarios necessitate the long-text generation ability of LLMs in these domains but provided with restricted amounts of usable data due to removing the identifying details for privacy. To expand the facilitation of clinical situations, we adopt LFQA benchmark datasets to explore how well open foundation LLMs respond to the content that consumers or patients typically inquire about, utilizing benchmarks that gather such inquiries (Singhal et al., 2023; Manes et al., 2024; Abacha et al., 2017; 2019).
@@ 2.2 Evaluating Long-Text Generation
The main challenge in conducting comprehensive research on the LFQA benchmark is the difficulty in automatic evaluation (Xu et al., 2023). Prior works provide various metrics for evaluating language models’ responses such as comparing the quality of machine-generated text to reference text (Lin, 2004; Ganesan, 2018) and capturing non-trivial semantic similarities (Papineni et al., 2002; Sellam et al., 2020; Zhang et al., 2019). With the increasing demand for using responses generated by LLMs, concurrent research also focuses on whether these responses accurately contain factual content and avoid generating false knowledge (i.e., hallucination) (Wei et al., 2024; Lee et al., 2022; Lin et al., 2022; Pal et al., 2023; Tian et al., 2023; Zhang et al., 2023; Kang et al., 2024; Lin et al., 2024; Dahl et al., 2024; Li et al., 2024a).
A widely known metric that can be used to measure factuality is FActScore (Min et al., 2023), which decomposes LLM responses into atomic facts and checks if they are supported by the source text. Additionally, there are metrics like Hallucination and Comprehensiveness (Manes et al., 2024) that measure the inclusion of crucial claims in the clinical domain. In detail, Hallucination (Manes et al., 2024) is a metric used to measure how many clinical claims are contradicted by the response of language models ($\hat{P}$). We compute the score as below,
@@
where $S$ refers to all statements containing Must Have (MH) and Nice to Have (NH) statements (i.e., $|S|=|MH|+|NH|$). Also, Comprehensiveness (Manes et al., 2024) is a metric used to measure how many clinically crucial claims are included in the response of language models. We compute the score as follows:
@@
To predict the entailment of the response, we use a classification model BioBERT (Lee et al., 2020) trained on NLI datasets (Bowman et al., 2015; Williams et al., 2018) on behalf of GPT-3.5-turbo due to the costs of API Calls. We provide detailed experiments in Appendix A.6. Also, we will describe the usage of these statements in the following section (Section 3.1). Our work is based on using these fine-grained and cost-effective evaluation metrics to understand how LLMs generate long-form text prioritizing factuality, semantic similarities, and word composition.
Figure 1: Current LFQA benchmark datasets lack comprehensive evaluation criteria, featuring just a pair of questions and answers (or not even an answer). In MedLFQA, we provide GPT-4 generated answers as well as two crucial statements to address this limitation. For instance, a well-generated GPT-4 response provides information on the definition, advantages, disadvantages, and side effects of Lexapro in response to a patient’s inquiry about it. Additionally, the answers and statements are structured to enable assessment of how closely the LLM response aligns with the correct answer in terms of multifaceted automatic evaluation: factuality, semantic similarity, and word composition.
@@ 3.1 Reconstruction of Biomedical Long-Form Question-Answering Datasets
The essential part of answering the patient’s question is conveying factual claims without false knowledge. To this end, the authors from (Manes et al., 2024) provide 1,212 patient questions originating from real-world conversations held on AI-driven clinical platform (i.e., K Health) containing long-form answers and two optional statements: Must Have Statements indicating that a model must include this statement to be medically accurate (e.g., providing all contraindications for a drug) and Nice to Have Statements indicating that the statements are supplemental (e.g., providing additional conditions where this drug may be helpful). These two statements provide an effective way to conduct an automatic evaluation of identifying factuality. Although the pairs of questions and answers are curated by medical experts, the dataset containing long-form answers is only limited to 202 pairs.
In this work, we introduce MedLFQA, which is constructed by expanding and reformulating current LFQA benchmark datasets to evaluate models’ responses automatically. To this end, we gather four biomedical LFQA datasets: LiveQA (Abacha et al., 2017), MedicationQA (Abacha et al., 2019), HealthSearchQA (Singhal et al., 2023), and K-QA (Manes et al., 2024). We describe the statistics of the benchmark datasets in Table 1. Each MedLFQA instance is comprised of four components: question (Q), long-form answer (A), Must Have statements (MH), and Nice to Have statements (NH). Specifically, LiveQA and MedicationQA datasets contain patients’ questions and their medical experts’ answers. HealthSearchQA only includes patients’ questions without their answers and crucial claims. In the K-QA dataset, the remaining examples (83%) that only consist of consumer questions are referred to as the K-QA Silver dataset.
In detail, if a medical expert’s answer exists, we create the two statements by decomposing the answer. For datasets containing only patients’ questions, we generate answers and statements using proprietary large language models such as GPT-4.333We provide detailed prompt in Appendix Table 12 For example, Figure 1 shows that the long-form answer generated by GPT-4 contains essential information, such as the pros and cons effects of Lexapro, compared to the golden answer that is curated with medical experts. We qualify the generated answers and statements through medical experts and provide the details in further section 3.2.
@@ 3.2 Qualification of Generated Answers and Statements
Our primary focus is to assess, through pairwise evaluation, whether GPT-4s’ answers are practically usable compared to thos of medical experts. Thus, we qualify the validity of predictions generated by GPT-4 using the K-QA golden dataset, whose answers are curated by medical experts. In order to assess a better response, we employ nine evaluation criteria from MedPALM: alignment with medical consensus (MC), reading comprehension (RC), knowledge recall (KC), reasoning (R), inclusion of irrelevant content (IRC), omission of important information (OII), potential for demographic bias (PDB), possible harm extent (PHE), possible harm likelihood (PHL). Using the criteria, we conduct a pairwise evaluation between GPT-4 predictions and K-QA golden answers through medical experts. [^4] Three medical experts, all at the resident level or higher, ensure that they were sufficiently qualified in medical knowledge. we recruit them from Seoul National University Hospital and have no conflict of interest. Additionally, we check an agreement by determining if at least two out of three medical experts choose the same answer. We want to note that our observation provide a high level of agreement among the experts across all criteria. [^5] We describe the details of these evaluation criteria in Appendix A.1
@@
We use GPT-4 to decompose answers and generate MH and NH statements, as described in K-QA dataset (Manes et al., 2024). According to the paper (Manes et al., 2024), a panel of six medical doctors, who were not involved in the initial decomposition of answers, utilized GPT-4 with few-shot prompting to generate these statements. They then curated the results by adding or removing statements, verifying only 6.86% of the automatically generated statements. This means that 93.14% of the statements produced by GPT-4 with few-shot prompting were left unchanged. Thus, we believe that if we could verify the answers generated for the patient questions are accurate, the statements derived from these answers are likely to be highly accurate as well.